OpenStudy (anonymous):

Maureen is taking an antibiotic. Her body metabolizes the antibiotic over a certain time interval. The table below shows the amount of antibiotic f(t), in mg, that was present in her body after time t: t (hours) 1 2 3 4 5 f(t) (mg) 150 90 54 32.4 19.4 Ken was administered 200 mg of the same antibiotic. The amount of antibiotic f(t) in his body after time t is shown by the equation below: f(t) = 200(0.976)t

OpenStudy (anonymous):

Which statement best describes the rate at which Maureen’s and Ken's bodies eliminated the antibiotic between 1 hour to 4 hours? Maureen's body eliminated the antibiotic faster than Ken's body. Maureen's body eliminated the antibiotic at the same rate as Ken's body. Maureen's body eliminated the antibiotic at half of the rate at which Ken's body eliminated the antibiotic. Maureen's body eliminated the antibiotic at one-fourth of the rate at which Ken's body eliminated the antibiotic.

OpenStudy (anonymous):

@ganeshie8 @jdoe0001 @dan815 @aaronq @Venny

OpenStudy (aaronq):

Is the t supposed to be an exponent, like \(\sf f(t) = 200(0.976)^t\) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i think so.

OpenStudy (aaronq):

Okay, so what the question asks is to compare the rates of clearance. Find the concentrations for Ken at t=1, 2, 3, 4 to compare them with Maureen's.

OpenStudy (anonymous):

really not understnding

OpenStudy (aaronq):

Find the amounts left at t=1, 2, 3, 4 Ken: \(\sf f(t)=200(0.976)^t\) \(\sf f(1)=195.2 ~mg\) \(\sf f(2)=190.5~mg\) etc

OpenStudy (anonymous):

is this what i compare

OpenStudy (aaronq):

You compare Ken's data (above) with Maureen's data (given in the question).

OpenStudy (anonymous):

its the same.

OpenStudy (anonymous):

Maureen's body eliminated the antibiotic at the same rate as Ken's body.

OpenStudy (aaronq):

thats not what it seems like to me Maureen from t=1 to t=2, went from 150 to 90 a loss of 40 mg in 1 hour Ken lost ~5mg in that same first hour

OpenStudy (anonymous):

so kens body eliminated it faster? how?

OpenStudy (anonymous):

What did you multiply?

OpenStudy (anonymous):

not ken maureen.

OpenStudy (aaronq):

yeah, look at the rates between the same time intervals

OpenStudy (anonymous):

Maureens body eliminated it faster.

OpenStudy (aaronq):

indeed!

OpenStudy (anonymous):

A

OpenStudy (anonymous):

can you help with 1 more

OpenStudy (aaronq):

sure thing!

OpenStudy (anonymous):

The functions f(x) and g(x) in the table below show Jane’s and Mariah's savings respectively, in dollars, after x days. Some values are missing in the table: x(years) 1 2 3 f(x) = 2x Jane’s savings in dollars 2 4 g(x) = 2x Mariah’s savings in dollars 2 4 Which statement best describes Jane and Mariah's savings in the long run? Jane will save more than Mariah. Mariah will save more than Jane. The savings will increase by the same factor for both. The savings will increase by the same percentage for both.

OpenStudy (anonymous):

I chose B

OpenStudy (aaronq):

Are both of the functions \(2x\) or \(2^x\)?

OpenStudy (anonymous):

its 2x and 2^x

OpenStudy (anonymous):

they are both different

OpenStudy (aaronq):

You can see that \(2x\) is linear and will increase at a constant rate, but \(2^x\) is exponential and will increase MUCH faster 

OpenStudy (anonymous):

so B.

OpenStudy (anonymous):

Mariah will save more than Jane.

OpenStudy (anonymous):

since its going up does that mean there's a loss in the money she has?

OpenStudy (aaronq):

nope, the graphs are their savings, no money is being lost here. The one with the exponential function will save more, but i dont know which on of the persons has it because they look the same to me

OpenStudy (anonymous):

what do you mean the same?

OpenStudy (anonymous):

The percentage?

OpenStudy (aaronq):

i mean from what you pasted, both of the functions look like "2x" so i dont know which one of the girls in the question has the exponential function.

OpenStudy (anonymous):

f(x) = 2x Jane’s savings in dollars

OpenStudy (anonymous):

g(x) = 2^x Mariah’s savings in dollars

OpenStudy (anonymous):

sorry i confused you

OpenStudy (aaronq):

no worries! so yeah you were right, Mariah will save more, so B would be right

OpenStudy (anonymous):

Thank so much :)

OpenStudy (aaronq):

no problem :)